OpenStudy (anonymous):

COMBINATIONS QUESTION: In how many ways can a committee of 6 boys & 6 girls be formed from a group of 10 boys & 15 girls?

OpenStudy (anonymous):

Probability = number of favorable ways / total number of ways We use the formula for combination in solving for the numerator and denominator. The combination of n things taken r at a time is C = n! / (n-r)!r! To solve for the total number of ways, we have C = 25! / (25-12)! 12! = 25! / 13! 12! = 5200300 This is the denominator. T solve for the numerator, boys x girls 10! / (10-6)! 6! x 15! / (15-6)! 6! 210 x 5005 = 1051050 This is the numerator. So we have 1051050 / 5200300 = 3003 / 14858 = 0.202

OpenStudy (anonymous):

How come you multiply and not add the two cases per gender?

OpenStudy (anonymous):

From what I understand, when the word used is 'and', you multiply it. If the word used is 'or', you add it.

OpenStudy (anonymous):

Ohhh, okay, thank you!